A chemical plant uses electrical energy to decompose aqueous solutions of NaCl to give Cl2, H2, and NaOH:

2NaCl(aq)+2H2O(l)→2NaOH(aq)+H2(g)+Cl2(g)

1. If the plant produces 1.6×106 kg (1600 metric tons) of Cl2 daily, estimate the quantities of H2 produced.

2. If the plant produces 1.6×106 kg (1600 metric tons) of Cl2 daily, estimate the quantities of NaOH produced.
Answer #1

1)

Molar mass of Cl2 = 70.9 g/mol

mass of Cl2 = 1.6*10^9 g

mol of Cl2 = (mass)/(molar mass)

= 1.6*10^9/70.9

= 2.257*10^7 mol

Balanced chemical equation is:

2NaCl(aq)+2H2O(l)→2NaOH(aq)+H2(g)+Cl2(g)

According to balanced equation

mol of H2 formed = moles of Cl2

= 2.257*10^7 mol

Molar mass of H2 = 2.016 g/mol

mass of H2 = number of mol * molar mass

= 2.257*10^7*2.016

= 4.55*10^7 g

= 4.55*10^4 Kg

Answer: 4.6*10^4 Kg

2)

According to balanced equation

mol of NaOH formed = (2/1)* moles of Cl2

= (2/1)*2.257*10^7

= 4.513*10^7 mol

Molar mass of NaOH,

MM = 1*MM(Na) + 1*MM(O) + 1*MM(H)

= 1*22.99 + 1*16.0 + 1*1.008

= 39.998 g/mol

mass of NaOH = number of mol * molar mass

= 4.513*10^7*40

= 1.805*10^9 g

= 1.805*10^6 Kg

Answer: 1.8*10^6 Kg